# Tony Lindeberg1    Zahra Babaiee2 and Peyman M. Kiasari2

### v1

@@ 3 Modelling the 8 “master key filters” in terms of idealized scale-space operations
Denoting the operators that extract the diagonal elements in the discrete weighted variance-based spatial spread measure $V_{T(\cdot)}(h(\cdot))$ in (4) by $V_{xx,T(\cdot)}(h(\cdot))$ and $V_{yy,T(\cdot)}(h(\cdot))$, respectively, we will in the following determine the scale estimates in idealized models of the “master key filters” according to (93)–(100), based on the criterion that the differences between the diagonal elements of the variance-based spatial spread measures should be as small as possible, between the idealized receptive field models and the learned filters, as specified in Equations (117)–(118) in Figure 6.
Filter $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$
1 0.644 0.583
2 0.586 0.644
3 0.690 0.674
4 0.756 0.460
5 1.107 0.945
6 0.900 0.889
8 0.609 0.601
Table 8: Estimates of the horizontal and vertical scale parameters $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ computed according to Method B, based on requiring the horizontal vs. the vertical discrete weighted variance-based spatial spread measures for the idealized models of the “master key filters” to be equal to the weighted discrete spatial spread measures for the the corresponding learned filters.
Table 8 shows the result of computing horizontal and vertical scale estimates $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{x,i}$ in this way for the 8 different “master key filters”.
Figure 7 shows visualizations of the corresponding idealized receptive field models of these filters.
Idealized 1(C1) Idealized 2(C1) Idealized 3(C1) Idealized 4(C1)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(C1) Idealized 6(C1) Idealized 7(C2) Idealized 8(C1)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method C1 in Section 3.3, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using different values of the scale parameters in the horizontal and the vertical directions according to (121). (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{x,1},\sigma_{x,1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{x,2},\sigma_{x,2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{x,3},\sigma_{x,3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{x,4},\sigma_{x,4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{x,5},\sigma_{x,5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{x,6},\sigma_{x,6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{x,8},\sigma_{x,8})$. (The numerical values of $\sigma_{i}$ for $i\in\{1,2,3,4,5,6,8\}$ are according to Table 9, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
Filter $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$
1 0.360 0.510
2 0.555 0.453
3 0.701 0.655
4 0.563 0.384
5 1.309 0.875
6 0.973 1.171
8 0.637 0.587
Table 9: Estimates of the horizontal and vertical scale parameters $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ computed according to Method C1, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(100) and the normalized learned filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73).
@@
Table 9 shows the resulting horizontal and vertical scale estimates obtained in this way. Figure 8 shows visualizations of the corresponding idealized receptive fields.
Filter $\hat{\sigma}_{i}$
1 0.458
2 0.448
3 0.671
4 0.420
5 1.387
6 1.090
8 0.611
Table 10: Estimates of the joint horizontal and vertical scale parameters $\hat{\sigma}_{i}$ computed according to Method C2, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(100) and the normalized learned filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73), under the assumption that the values of the horizontal and the vertical scale parameters are the same $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$.
Idealized 1(C2) Idealized 2(C2) Idealized 3(C2) Idealized 4(C2)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(C2) Idealized 6(C2) Idealized 7(C2) Idealized 8(C2)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method C2 in Section 3.4, based on minimizing the discrete $l_{1}$-norm between the idealized receptive field models and the normalized versions of the learned filters according to (120). (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{1},\sigma_{1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{2},\sigma_{2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{3},\sigma_{3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{4},\sigma_{4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{5},\sigma_{5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{6},\sigma_{6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{8},\sigma_{8})$. (The numerical values of $\sigma_{i}$ for $i\in\{1,2,3,4,5,6,8\}$ are according to Table 10, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
@@
As a complement to the above analysis, let us also consider the special case when scale parameters are required to be the same in the horizontal and the vertical directions $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$, for which we then determine each scale estimate $\hat{\sigma}_{i}$ from the criterion
@@
Table 10 shows the result of computing scale estimates for the idealized receptive field models in this way. Figure 9 shows visualizations of the resulting receptive fields.
Idealized 1(D1) Idealized 2(D1) Idealized 3(D1) Idealized 4(D2)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(D1) Idealized 6(D1) Idealized 7(D2) Idealized 8(D1)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method D1 in Section 3.5, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using different values of the scale parameters in the horizontal and the vertical directions according to (121). (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{x,1},\sigma_{x,1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{x,2},\sigma_{x,2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{x,3},\sigma_{x,3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{x,4},\sigma_{x,4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{x,5},\sigma_{x,5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{x,6},\sigma_{x,6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{x,8},\sigma_{x,8})$. (The numerical values of $\sigma_{i}$ for $i\in\{1,2,3,4,5,6,8\}$ are according to Table 11, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
Filter $\hat{\sigma}_{x,i}$ $\hat{\sigma}_{y,i}$
1 0.491 0.722
2 0.581 0.519
3 0.483 0.503
4 0.500 0.000
5 1.300 1.004
6 0.984 1.074
8 0.615 0.608
Table 11: Estimates of the horizontal and vertical scale parameters $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{y,i}$ computed according to Method D1, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(100) and the normalized learned filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73).
@@
In analogy with the above $l_{1}$-norm-based method in Section 3.3, for determining separate scale parameters in the horizontal and the vertical directions for idealized receptive field models, we can also an $l_{2}$-norm-based method from the criterion
@@
Table 11 shows the horizontal and the vertical scale estimates obtained in this way. Figure 10 shows visualizations of the corresponding idealized receptive fields.
Filter $\hat{\sigma}_{i}$
1 0.644
2 0.558
3 0.495
4 0.380
5 1.193
6 1.038
8 0.612
Table 12: Estimates of the joint horizontal and vertical scale parameters $\hat{\sigma}_{i}$ computed according to Method D2, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field model $h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{x,1},\sigma_{y,1})$ according to (93)–(100) and the normalized learned filter $h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)$ according to (63)–(68) and (73), under the assumption that the values of the horizontal and the vertical scale parameters are the same $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$.
Idealized 1(D2) Idealized 2(D2) Idealized 3(D2) Idealized 4(D2)
Refer to caption Refer to caption Refer to caption Refer to caption
Idealized 5(D2) Idealized 6(D2) Idealized 7(D2) Idealized 8(D2)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Idealized models of the 8 “master key filters” in Figure 2, in terms of combinations pure difference operators with spatial smoothing filters based on the discrete analogue of the Gaussian kernel, with the amount of spatial smoothing determined from the spatial extents of the 8 “master key filters” using Method D2 in Section 3.6, based on minimizing the discrete $l_{2}$-norm between the idealized receptive field models and the normalized versions of the learned filters and using the same values of the scale parameters in the horizontal and the vertical directions according to (120). (Idealized 1): Non-centered vertical difference operator $(\delta_{y+}T)(m,n;\;\sigma_{1},\sigma_{1})$. (Idealized 2): Non-centered horizontal difference operator $(\delta_{x-}T)(m,n;\;\sigma_{2},\sigma_{2})$. (Idealized 3): Non-centered vertical difference operator $(\delta_{y-}T)(m,n;\;\sigma_{3},\sigma_{3})$. (Idealized 4): Non-centered horizontal difference operator $(\delta_{x+}T)(m,n;\;\sigma_{4},\sigma_{4})$. (Idealized 5): Centered horizontal difference operator $(\delta_{x}T)(m,n;\;\sigma_{5},\sigma_{5})$. (Idealized 6): Centered vertical difference operator $(\delta_{y}T)(m,n;\;\sigma_{6},\sigma_{6})$. (Idealized 7): Sharpening operation $1-\gamma_{7}\,(\nabla_{5}^{2}T)(m,n;\;\sigma_{7},\sigma_{7})$. (Idealized 8): Pure spatial smoothing $T(m,n;\;\sigma_{8},\sigma_{8})$. (The numerical values of $\sigma_{i}$ for $i\in\{1,2,3,4,5,6,8\}$ are according to Table 12, and with $\sigma_{7}$ additionally according to Table 13.) (Note that the contrasts of Filters 2 and 3 are reversed in relation to the sign conventions for corresponding “master key filters”.)
@@
As a complement to the above analysis, let us also consider the special case when scale parameters are required to be the same in the horizontal and the vertical directions $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$, for which we then determine each scale estimate $\hat{\sigma}_{i}$ from the criterion
$\hat{\sigma}_{i}=\operatorname{argmin}_{\hat{\sigma}_{i}}\|\Delta h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot;\;\sigma_{i},\sigma_{i})\|_{2}=\\ =\operatorname{argmin}_{\sigma_{i}}\|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{i},\sigma_{i})-h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)\|_{2}.$ (122)
Table 12 shows the result of computing scale estimates for the idealized receptive field models in this way. Figure 11 shows visualizations of the resulting idealized receptive fields.
@@ 3.7 Two-parameter modelling of Filter 7 as a sharpening filter
Visualizations of the corresponding filter maps are given among the visualizations of the other filter maps for Method C2 and Method D2 in Figures 9 and 11, respectively.
Method A Method B Method C1 Method C2 Method D1 Method D2
1 0.580 0.558 0.644 0.583 0.360 0.510 0.458 0.491 0.722 0.644
2 0.559 0.580 0.586 0.644 0.555 0.453 0.448 0.581 0.519 0.558
3 0.617 0.601 0.690 0.674 0.701 0.655 0.671 0.483 0.503 0.495
4 0.642 0.424 0.756 0.460 0.563 0.384 0.420 0.500 0.000 0.380
5 1.075 0.800 1.107 0.945 1.309 0.875 1.387 1.300 1.004 1.193
6 0.771 1.003 0.900 0.889 0.973 1.171 1.090 0.984 1.074 1.038
7 0.654 0.675
8 0.552 0.545 0.609 0.601 0.637 0.587 0.611 0.615 0.608 0.612
Table 14: Summary of the scale estimates obtained for the 8 “master key filters” using the different criteria using for computing them in different ways using: (i) Method A based on direct lookup of anisotropic scale estimates from variance-based spatial spread measures, (ii) Method B based on anisotropic matching of discrete spatial spread measures between the idealized receptive field models and the learned filters, (iii) Method C1 based on anisotropic discrete $l_{1}$-norm minimization, (iv) Method C2 based on isotropic discrete $l_{1}$-norm minimization, (v) Method D1 based on anisotropic discrete $l_{2}$-norm minimization and (vi) Method D2 based on isotropic discrete $l_{1}$-norm minimization. These methods have been explored for all the learned filters except for Filter 7, for which instead the two-parameter modelling scheme in Section 3.7, which also determines a scaling constant for the negative Laplacian-of-the-Gaussian filter in the sharpening operation.
Table 14 gives a summarizing overview of the scale estimates obtained from the different methods applied to the individual learned filters.
@@
A notable property of the learned receptive fields is that their orientational preferences are very closely aligned to the directions of Cartesian coordinate directions. In our preliminary investigations in Section 2.2, we noted that the unweighted spatial covariance matrices of the absolute values of the learned filter kernels were very close to diagonal (see Table 1). In our follow-up analysis in Section 2.5, we also noted that the analogous weighted spatial covariance matrices of the absolute values of the learned filter kernels were also very close to diagonal (see Table 6). This overall property then motivated us to state separable idealized models of the learned receptive fields according to (93)–(100), which we then in Section 3 demonstrated to provide qualitatively reasonable models of the shapes of the learned filters.
Given that natural image structures usually show a substantial variability with regard to orientations in the image domain, one may then ask how such a variability could be spanned by the deep network in which the learned filters constitute the primary computational primitives. From the property of the first-order directional derivative operator $\partial_{\varphi}$ in the direction $\varphi$
@@
it does, however, follow that the computation of derivatives in the horizontal and vertical directions, as performed by the operators $\partial_{x}$ and $\partial_{y}$, can be linear combinations span the space of all directions $\varphi$ in the 2-D image plane. Thus, since shapes of Filters 1–6 can be qualitatively well described by first-order Gaussian derivative kernels in the Cartesian coordinate directions, we could regard the depthwise combinations of the responses from the learned filters as having the potential ability to span a reasonable approximation of the space of the first-order directional derivatives in any direction $\varphi$, at two separate scales: both (i) the joint scale of Filters 1–4 and (ii) the joint scale of Filters 5–6.
Secondly, one may ask why there is a clear preference to first-order spatial derivatives in the learned “master key filters”, and not regarding e.g. second-order spatial derivatives, and why the receptive fields of Filters 1–4 are not fully centered, but instead shifted by approximately half a grid spacing along its preferred orientation in the image plane. From properties of the non-centered primitive first-order difference operators $\delta_{x+}$, $\delta_{x-}$, $\delta_{y+}$ and $\delta_{y-}$, we can, however, first of all by linear combinations construct their corresponding centered first-order difference operators according to
$\begin{split}\delta_{x}=\frac{\delta_{x+}+\delta_{x-}}{2},\end{split}$ (126)
$\begin{split}\delta_{y}=\frac{\delta_{y+}+\delta_{y-}}{2}.\end{split}$ (127)
@@
$\begin{split}\delta_{xx}=\delta_{x+}-\delta_{x-},\end{split}$ (128)
$\begin{split}\delta_{yy}=\delta_{y+}-\delta_{y-}.\end{split}$ (129)
Thus, by linear combinations of the output from Filters 1–4, it should be possibly to to a certain degree of approximation span the space of second-order directional derivatives along the Cartesian coordinate directions at the joint scale of Filters 1–4.
@@
Given this theoretical motivation, one could thus raise the question of whether the performance of depthwise-separable deep networks with the receptive field shapes replaced by idealized receptive field models could be improved by extending the set of “master key filters” with a filter that constitutes a discrete approximation of the mixed second-order Gaussian derivative operator at the joint scale of Filters 1–4:
@@
Concerning these interpretations, it should finally be emphasized that in depthwise-separable CNN:s, no direct linear combinations of the depthwise filters are performed. The depthwise convolutions are applied on distinct input feature maps, and the following pointwise layers linearly combine the outputs of these convolutions. Hence, the interpretations of the receptive field spanning different subspaces should be interpreted in terms of gross properties at the level of the population of the receptive fields only, and not concerning individual receptive fields.

### v2

@@ 4 Modelling the 8 “master key filters” in terms of idealized scale-space operations
Let us denote the operators that extract the diagonal elements in the discrete weighted variance-based spatial spread measure $V_{T(\cdot)}(h(\cdot))$ in (4) by $V_{xx,T(\cdot)}(h(\cdot))$ and $V_{yy,T(\cdot)}(h(\cdot))$, respectively. In the following, we will determine the scale estimates in idealized models of the “master key filters” according to (95)–(102), based on the criterion that the differences between the diagonal elements of the variance-based spatial spread measures should be as small as possible, between the idealized receptive field models and the learned filters, as specified in Equations (112)–(113) in Figure 6.
The second major column in Table 7 shows the result of computing horizontal and vertical scale estimates $\hat{\sigma}_{x,i}$ and $\hat{\sigma}_{x,i}$ in this way for the 8 different “master key filters”.
The 3rd row in Figure 2 shows visualizations of the corresponding idealized receptive field models of these filters.
@@
The third major column in Table 7 shows the resulting horizontal and vertical scale estimates obtained in this way. The 4th row in Figure 2 shows visualizations of the corresponding idealized receptive fields.
@@
As a complement to the above analysis, let us also consider the special case when the scale parameters are required to be the same in the horizontal and the vertical directions $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$, for which we then determine each scale estimate $\hat{\sigma}_{i}$ from the criterion
@@
The fourth major column in Table 7 shows the result of computing scale estimates for the idealized receptive field models in this way. The 5th row in Figure 2 shows visualizations of the resulting receptive fields.
@@
In analogy with the above $l_{1}$-norm-based method in Section 4.3, for determining separate scale parameters in the horizontal and the vertical directions for idealized receptive field models, we can also use an $l_{2}$-norm-based method from the criterion
@@
The fifth major column in Table 7 shows the horizontal and the vertical scale estimates obtained in this way. The 6th row in Figure 2 shows visualizations of the corresponding idealized receptive fields.
@@
Similarly to the case with $l_{1}$-norm minimization, we can also require the scale parameters in the horizontal and the vertical directions $\sigma_{x,i}=\sigma_{y,i}=\sigma_{i}$ to be equal, and determine each scale estimate $\hat{\sigma}_{i}$ from the criterion
$\hat{\sigma}_{i}=\operatorname{argmin}_{\hat{\sigma}_{i}}\|\Delta h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot;\;\sigma_{i},\sigma_{i})\|_{2}=\\ =\operatorname{argmin}_{\sigma_{i}}\|h_{i,\scriptsize\mbox{ideal}}(\cdot,\cdot;\;\sigma_{i},\sigma_{i})-h_{i,\scriptsize\mbox{norm}}(\cdot,\cdot)\|_{2},$ (117)
see the sixth major column in Table 7 for resulting the resulting scale estimates for the idealized receptive field models, and the 7th row in Figure 2 for visualizations of the resulting idealized receptive fields.
@@ 4.7 Two-parameter modelling of Filter 7 as a sharpening filter
Visualizations of the corresponding filter maps are given among the visualizations of the other filter maps for Method C2 and Method D2 in the 5th and 7th rows in Figure 2, respectively.
@@
A notable property of the learned receptive fields is that their orientational preferences are very closely aligned to the directions of Cartesian coordinate directions. In our preliminary investigations in Section 3.2, we noted that the unweighted spatial covariance matrices of the absolute values of the learned filter kernels were very close to diagonal (see Table 1). In our follow-up analysis in Section 3.5, we also noted that the analogous weighted spatial covariance matrices of the absolute values of the learned filter kernels were also very close to diagonal (see Table 6). This overall property then motivated us to state separable idealized models of the learned receptive fields according to (95)–(102), which we demonstrated in Section 4 to provide qualitatively reasonable models of the shapes of the learned filters.
Given that natural image structures usually show a substantial variability with regard to orientations in the image domain, one may then ask how such a variability could be spanned by the deep network, in which the learned filters constitute the primary computational primitives. From the property of the first-order directional derivative operator $\partial_{\varphi}$ in the direction $\varphi$
@@
it does, however, follow that the computation of derivatives in the horizontal and vertical directions, as performed by the operators $\partial_{x}$ and $\partial_{y}$, can by linear combinations span the space of all directions $\varphi$ in the 2-D image plane. Thus, since shapes of Filters 1–6 can be qualitatively well described by first-order Gaussian derivative kernels along the Cartesian coordinate directions, we could regard the depthwise combinations of the responses from the learned filters as having the potential ability to span a reasonable approximation of the space of the first-order directional derivatives in any direction $\varphi$, at two separate scales: both (i) at the joint scale of Filters 1–4 and (ii) at the joint scale of Filters 5–6.
Secondly, one may ask why there is a clear preference to first-order spatial derivatives in the learned “master key filters”, and not regarding e.g. second-order spatial derivatives, and why the receptive fields of Filters 1–4 are not fully centered, but instead shifted by approximately half a grid spacing along its preferred orientation in the image plane. From properties of the non-centered primitive first-order difference operators $\delta_{x+}$, $\delta_{x-}$, $\delta_{y+}$ and $\delta_{y-}$, we can, however, by linear combinations construct their corresponding centered first-order difference operators according to
$\delta_{x}$ $=\frac{\delta_{x+}+\delta_{x-}}{2},$ $\delta_{y}$ $=\frac{\delta_{y+}+\delta_{y-}}{2}.$ (121)
@@
$\delta_{xx}$ $=\delta_{x+}-\delta_{x-},$ $\delta_{yy}$ $=\delta_{y+}-\delta_{y-}.$ (122)
Thus, by linear combinations of the output from Filters 1–4, it should be possible to a certain degree of approximation to span the space of second-order directional derivatives along the Cartesian coordinate directions at the joint scale of Filters 1–4.
@@
Given this theoretical motivation, one could thus raise the question about whether the performance of depthwise-separable deep networks with the receptive field shapes replaced by idealized receptive field models could be improved by extending the set of “master key filters” with a filter that constitutes a discrete approximation of the mixed second-order Gaussian derivative operator at the joint scale of Filters 1–4:
@@
Concerning these interpretations, it should finally be emphasized that in depthwise-separable CNNs, no direct linear combinations of the depthwise filters are performed. The depthwise convolutions are applied on distinct input feature maps, and the following pointwise layers linearly combine the outputs of these convolutions. Hence, the interpretations of the receptive field spanning different subspaces should be interpreted in terms of gross properties at the level of the population of the receptive fields only, and not concerning individual receptive fields.